OpenStudy (anonymous):

Solve for x. x/4+ 3/4=6

OpenStudy (anonymous):

Use order of operation backward for solving these equations. Please Excuse My Dear Aunt Sally Parenthese Exponent Multiplication Division Addition Subtraction So first subtract 3/4 from both sides of the equation. Then multiply both sides by 4 to get x. What would be your final answer?

OpenStudy (anonymous):

How do I subtract 3/4 from each side?

OpenStudy (anonymous):

Do you know how to solve 6 - (3/4)?

OpenStudy (anonymous):

No I'm completing a study guide this is all new to me. Could you show me how to solve this one answer so I can try the others on my own?

OpenStudy (ghazi):

x/4=6-3/4=(24-3)/4...just take LCM then x= (4*21)/4=21

OpenStudy (anonymous):

Thank youuu :)

OpenStudy (ghazi):

:)